# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

### v4

@@ Abstract
Peer effects significantly shape behavior, attitudes, and performance, particularly in elementary and middle school settings where younger students are highly susceptible to peer influences. By enhancing peer effects, this study uses data from the China Educational Panel Survey (CEPS) to design a classroom assignment policy that maximizes educational outcomes. Our approach comprises three steps: firstly, we develop a friendship formation discrete choice model and estimate it with an interpretable neural network architecture, PeerNN, generating an adjacency-probability matrix $\Omega$ that reflects friendship formation probabilities. In our empirical study, $\Omega$ aligns with well-known social network characteristics such as gender homophily, the presence of central nodes, and varying popularity levels across student subgroups. We also compare the out-of-sample prediction performance of PeerNN and the linear-in-means model. The former outperforms by a huge margin. Secondly, we incorporate $\Omega$ into a linear-in-means model to estimate peer effects. The peer effect parameter, $\beta$, has a different interpretation from the conventional linear-in-means model and opens up a strategic scope of mean-maximizing classroom assignment policy. By exploiting the conditional random classroom assignment in many Chinese middle schools, we construct a valid instrument to address the endogeneity issue induced by $\Omega$ and consistently estimate $\beta$. Lastly, utilizing the estimates of $\Omega$ and $\beta$, we employ a genetic algorithm (GA) to search for the mean-maximizing class assignment policy. Though the result is much more efficient (i.e. more positive average peer effect) than random classroom assignment (i.e. the current practice in most Chinese middle schools), GA policy is highly inequitable: a small number of students are predicted to experience severely negative peer effects. To balance students’ academic performance with educational equity, we propose a fairness metric and penalize classroom assignment that generates large variances in peer effects. The modified method is called algorithmically fair genetic algorithm (AFGA). AFGA policy is less efficient but much more equitable. We allow user-defined parameters for AFGA such that the school principals can adjust the trade-off between efficiency and equity according to their preferences. This comprehensive three-step framework not only improves class assignments but also offers a solution extendable to other group assignment challenges where peer effects are a consideration.
@@ 1 Introduction
Peer effects denote the impact individuals within the same social circles exert on each other, influencing various aspects of behavior, attitudes, and performance. Numerous studies (Feld and Zölitz, 2017; Carrell et al., 2018; Shure, 2021; Huang et al., 2021) have explored peer effects, particularly emphasizing their importance in classroom settings. Within the literature on classroom peer effects, one of the popular regression specifications is the linear-in-means model which assumes homogeneity in spillover effects among cohorts. Though the linear-in-means model is easy to implement and has a clear interpretation, it suffers from two major drawbacks. Firstly, the linear-in-means model assumes a uniform strength of the connection between any pair of individuals. Consequently, it does not shed light on network formation and can hardly capture differential peer effects individuals receive from their cohorts. Though under assumptions specified by Manski (1993), the linear-in-means model successfully identifies some parameters related to the average peer effect, researchers may be interested in the entire distribution of peer effect, or at least some aspects of the distribution that are not fully pinned down by the average. To uncover the distribution of peer effect, consideration of network formation is necessary. Linear-in-means model ignores intricacies of network development and therefore, struggles to provide insights on the distribution of peer effect. Secondly, neglecting the network/friendship development hinders the formulation of effective policies that can address issues arising from peer effects. This deficiency in capturing the nuances of network dynamics poses a challenge when optimizing group or class assignments. Therefore, exploring alternative models that account for the complexities of network formation is imperative for more comprehensive learning about peer effect and more effective design of group assignment policy.
In this paper, we expand the linear-in-means methodology by incorporating friendship formation into our model. The modification changes the interpretation of the peer effect parameter and opens up a strategic scope of classroom assignment optimization. Our methodology breaks a school principal’s classroom assignment problem into three parts: predicting friendship formation, measuring peer effect, and formulating an optimal class assignment policy informed by the outcomes of the first two components. Next, we elaborate further on the three-step empirical framework.
@@
First, we predict counterfactual friendship formations under different classroom assignment policies using a content-based approach rather than the conventional graph-based methods. In other words, we exclusively use students’ individual information as predictors, but not their network information as predictors. This approach is necessitated by the absence of linkage information under varying assignments, setting our work apart from mainstream link prediction studies in machine learning (Ahmed and Chen, 2016; Wu et al., 2016; Wang et al., 2014; Dhelim et al., 2022; Chunaev, 2020; Su et al., 2022; Samanta et al., 2021; Ling et al., 2023). Though both our work and these cited papers use neural networks to approximate complicated social network formation processes, we intentionally avoid techniques such as aggregating neighbors’ information which is commonly used in the cited papers. In contrast, we model friendship preferences using a micro-founded discrete choice setup and estimate it with PeerNN, a novel and interpretable neural network design. PeerNN captures social network characteristics such as gender homophily, outperforming traditional linear-in-means models used in peer effect literature (Guryan et al., 2009; Carrell and Hoekstra, 2010; Carrell et al., 2013; Burke and Sass, 2013). Another advantage of PeerNN is that it does not need expensive linkage data as a response variable. Instead, we can use easy-to-collect aggregate relational data (McCormick and Zheng, 2015; Breza et al., 2020; McCormick, 2020).
@@
Second, using the friendship prediction results from PeerNN, we measure peer effects. We replace the restrictive uniform spillover assumption of the linear-in-means model with a more reasonable friendship-weighted specification. This specification leverages the nonuniform friendship formation predicted by PeerNN, addressing the limitations of the linear-in-means model which assumes uniform spillover effects and ignores friendship dynamics (Bramoullé et al., 2009; Goldsmith-Pinkham and Imbens, 2013). However, network formation is endogenous. We exploit the random classroom assignment conditional on school choice to construct an instrument to conduct causal inference. The linear-in-means model measures the causal relationship between classmates’ average characteristics and a student’s outcome, whereas our specification quantifies the causal impact of friends’ weighted average characteristics on a student’s outcome. The change in interpretation of the causal parameter gives rise to a strategic scope for classroom assignment.
Note that there are myriads of channels peer effect can operate through. We focus our study on one specific channel through which peer effects can manifest, namely, the influence of a student’s friends’ average 6th-grade class quantile on the student’s 8th-grade cognitive ability (similar to Murphy and Weinhardt (2020)). The selection of this particular channel is not a central aspect of our research and researchers can readily adapt our framework to investigate alternative channels of peer effect.
@@
In the third step, we optimize classroom assignments in terms of maximizing peer effects. Unlike existing literature, which assumes straightforward dependence between network and group assignment policy (Bhattacharya, 2009; Graham et al., 2014), our problem allows for more intricate dependence between friendship formation and classroom assignment policy. We employ a heuristic optimization method: genetic algorithm (GA) to navigate this complex problem. GA outputs a classroom assignment policy that on average improves peer effect by 1.9% in our counterfactual simulation. However, the output policy is extremely inequitable. In pursuing the highest average peer effect for the whole classroom, GA gives up on a few disruptive peers by predicting that they form an exclusive clique. Consequently, students in that clique experience severely negative peer effects, but overall the entire classroom has a higher average peer effect. Such educational outcomes are extremely inequitable. We propose a modified fairness-aware fitness function for GA and name it algorithmically fair genetic algorithm (AFGA). The new fitness function penalizes the variance of predicted peer effects within and across classrooms. AFGA outputs slightly less efficient (the improvement is 1.2%) but much more equitable policies. This approach aligns with the work of Kleinberg et al. (2018) and Mitchell et al. (2021), who discuss the integration of algorithmic fairness in optimization methods.
This paper has six contributions. (1) We build a micro-founded discrete choice model for friendship formation, the model can be flexibly parameterized and estimated by the neural network method. Our model shows how flexible machine learning algorithms can be adapted to model complicated social processes while maintaining interpretability. The interpretability offers three advantages: (i) we can obtain a friendship probability-adjacency matrix, $\Omega$, as an intermediate product of PeerNN which is an integral part of peer effect estimation and classroom assignment optimization, (ii) the interpretability of $\Omega$ guides the design of loss function of PeerNN which aligns with the discrete choice model even when the response variable is aggregated relational and true linkages are not observed (iii) we can avoid tuning hyperparameters by examining the interpretability of $\Omega$. (2) We modify the classical linear-in-means model by replacing the uniform spillover assumption with a more reasonable friendship-weighted specification. The modified peer effect estimate is interpreted as how much influence a friend has on a student, rather than how much influence a classmate has on a student. The new interpretation opens up a strategic scope of classroom assignment policy optimization whereas the classical linear-in-means fails to do so. (3) Prior research has attempted to conduct causal inference under networked interference (Ma and Tresp, 2021; Jiang and Sun, 2022; Huang et al., 2023), which is exactly the setup in the classroom peer effect. Recognizing that in our data many classrooms are randomly assigned, we disentangle the network interference (counterfactual friendship prediction) and causal inference (peer effect measurement) problems. By exploiting the random assignment, we construct a valid instrument to isolate the peer effect from endogenous friendship selection. (4) Our three-step framework outputs an implementable classroom assignment policy. To the best of our knowledge, this is the first attempt to optimize classroom assignments with real data and consideration of friendship formation. Theoretical works (Berthelon et al., 2019; Bizzotto and Vigier, 2022; Sarkisian and Yamashita, 2024) do not bring their models to real data, whereas empirical works (Bhattacharya, 2009; Graham et al., 2014; Carrell et al., 2013) implicitly assume straightforward link formation like the linear-in-means model (though the peer effect parameter can be heterogeneous) which allows them to solve the optimization problem as mixed integer programming problems (MIPs). Our friendship model is more sophisticated, making the optimization problem not possible to be recast as an MIP. Therefore, we resort to a heuristic optimization method: genetic algorithm (GA). On top of adopting a novel approach (i.e. GA) to the optimization problem, we incorporate common institutional constraints such as gender ratio and classroom size into GA, tailoring it to the group assignment context. (5) While maximizing the average peer effect seems a natural objective of the classroom assignment problem, the plugged-in GA method outputs a policy that generates extremely unfair educational outcomes according to our counterfactual analysis. Therefore, we propose a fairness measure and penalize unfair educational outcomes to balance efficiency and equity. (6) Combining these three elements: predicting friendship formation, measuring peer effects, and optimizing classroom assignments, constitutes a comprehensive framework that extends beyond classroom setups. This framework applies to various group assignment contexts where peer effects are crucial, such as workplace teams (Anderson et al., 2008; Bandiera et al., 2010; Mas and Moretti, 2009), healthcare interventions (Pianese and Belfiore, 2021; Pagkas-Bather et al., 2020), and community programs (Valente et al., 2003; Yu et al., 2024; Cero et al., 2024).
@@
The rest of the paper is organized as follows: Section 2 relates our work to cutting-edge works in various research fields, Section 3 describes the data and introduces the notations that we use in this paper, Section 4 sets up the models and objective functions in our three-step framework, Section 5 demonstrates and explains the results that we obtain from the empirical study and Section 6 concludes the paper and discusses possible future research directions.
2 Literature review
In this section, we summarize how the problems we aim to tackle and the methods we use in this paper relate to and differ from existing literature.
2.1 Classroom assignment
This paper focuses on a specific channel of peer effect: how much a student’s friends’ average 6th-grade class quantile influences the student’s 8th-grade cognitive ability (similar to Murphy and Weinhardt (2020)). We address a principal’s classroom assignment problem: designing a policy rule to optimize the specified peer effect. The closest studies to our paper are Carrell et al. (2013) and Griffith (2024).
In the former paper, the authors conduct an experiment of class assignment aiming to maximize the positive influence of the high-ability students on the bottom 33% students’ GPA. In their study, the experiment result turns out to be surprising as the designed class assignment policy does not improve the bottom 33% students’ performance. The authors reconcile this counterintuitive result by arguing that the linear-in-means model does not consider dynamic friendship formation which aligns with our critique of the linear-in-means model. The first component of our three-step framework addresses this exact problem. We build a micro-founded discrete choice model to capture the friendship formation pattern which is then incorporated into our peer effect estimation and classroom assignment optimization.
The latter paper proves identification results for friendship formation and peer effect measurement, which this paper does not cover. However, to obtain the theoretical results, Griffith (2024) imposes much more restrictive parametric assumptions for friendship formation than PeerNN. Another advantage of PeerNN is that it does not require linkage data like the latter paper, making our paper more broadly applicable and less ethically concerning. In addition, our policy optimization strategy is more systematic. Griffith (2024) simulate a few counterfactual scenarios to compare a small number of assignment policies, our paper adapts GA to search for the optimal policy and subsequently considers algorithmic fairness which will be discussed in Section 2.7. The biggest difference between our paper and Griffith (2024) is that we aim to optimize group assignment policy and the network is policy-dependent, in contrast, the latter paper considers treatment assignment (same as papers cited in Section 2.6) and the network is fixed.
Other attempts to tackle classroom assignment problems or similar group assignment problems include Bhattacharya (2009), Graham et al. (2014), Berthelon et al. (2019), Bizzotto and Vigier (2022) and Sarkisian and Yamashita (2024), we defer the discussion of the difference between these works and ours to Section 2.8 when we compare their choices of optimization method to ours.
2.2 Link prediction for social network using neural networks
The friendship prediction problem in our work differs from the mainstream link prediction problems in social networks (later referred to as the mainstream problem). Daud et al. (2020) summarize that the mainstream problem focuses on (1) predicting future and missing links (Ahmed and Chen, 2016; Wu et al., 2016), (2) recommendation (Wang et al., 2014; Dhelim et al., 2022), (3) detection (Chunaev, 2020; Su et al., 2022), and (4) influence analysis (Samanta et al., 2021; Ling et al., 2023). In contrast, our work predicts counterfactual friendship formation under different classroom assignment policies. Unlike conventional approaches that utilize graph-based measures, we exclusively rely on content-based measures due to the absence of linkage information under different classroom assignments. For the same reason, common methods for analyzing network data, like aggregating neighbors’ information, are inapplicable to our setup. We use example C.1 to illustrate the key differences between our counterfactual friendship prediction problem and the mainstream problem. All examples are collected in Appendix C.
2.3 Modelling friendship formation with aggregate relational data
Our friendship prediction problem entails a counterfactual scenario, wherein we aim to understand how students would select friends under an alternative classroom assignment, operating on the assumption that friendships exclusively develop within the confines of the classroom. Due to the distinctive nature of our research, we are particularly interested in modeling friendship preference. As such, we build a micro-founded economic model that captures friendship preference heterogeneity. Dong et al. (2012); Zhang et al. (2014) adopt a similar concept when building recommendation systems across heterogeneous networks. We estimate the model with a novel and interpretable neural network design: PeerNN. It takes student characteristics as input and one of its intermediate outputs is an adjacency-probability matrix, $\Omega$. Based on our training results, the $\Omega$ matrix modeled by PeerNN captures commonly known social network characteristics including gender homophily, presence of central nodes, and heterogeneous levels of student popularity across distinct subgroups. In terms of network prediction, PeerNN substantially outperforms the linear-in-means model which is the predominant methodology in peer effect literature. Additionally, PeerNN does not require expensive friendship linkage data, instead, its response variable can be easy-to-collect aggregate relational data (McCormick, 2020; Breza et al., 2020; Laga et al., 2021).
2.4 Homophily and transitivity effect in sociology and economics
McPherson et al. (2001); Kossinets and Watts (2009) establish the importance of the homophily effect in network formation. This finding in sociology inspires a large amount of structural work in economics including McCormick and Zheng (2015); Currarini et al. (2016); Breza et al. (2020); Mele (2022). A related concept termed transitivity (i.e. one is more likely to befriend a friend’s friend) has also sparked many structural works in both sociology and economics including Richters and Peixoto (2011); Block (2015); Vasques Filho and O’Neale (2020). By incorporating these two concepts into the loss function of PeerNN, we estimate a structural model for friendship formation.
2.5 Measuring peer effect with linear-in-means model
Traditional studies in peer effect literature, including works by (Guryan et al., 2009; Carrell and Hoekstra, 2010; Carrell et al., 2013; Burke and Sass, 2013), have largely adopted the linear-in-means model, assuming uniform spillover effects among cohorts (See Appendix E for detail of linear-in-means model). This approach, while straightforward and interpretable (Bramoullé et al., 2009; Goldsmith-Pinkham and Imbens, 2013), overlooks the complexity of social networks and fails to capture the varied peer effects individuals experience (Angrist, 2014). By incorporating $\Omega$ into the standard linear-in-means framework, we relax this key assumption prevalent in previous peer effect studies to quantify a peer effect parameter with heightened policy relevance utilizing an instrumental variable approach. The statistically significant and positive peer effect estimate indicates the importance of maximizing the influence of high-achieving students from elementary school to enhance peer effects in middle school classrooms.
2.6 Causal inference under networked interference using neural network
PeerNN + IV + AFGA GNNCE NetEst SPNet
Prediction target Optimal policy Optimal policy Treatment effect Treatment effect
Observable confounder Instrumental variable DDP Adversarial learning DDP
Unobservable confounder Instrumental variable $\times$ $\times$ DDP + treatment prediction
Policy optimization Genetic algorithm Neural network $\times$ $\times$
Table 1: We propose a framework that combines neural network, instrumental variable, and genetic algorithm (PeerNN + IV + AFGA) to solve for optimal class assignment policy. Ma and Tresp (2021); Jiang and Sun (2022); Huang et al. (2023) focus on estimating causal inference using graph neural network (GNN). Their solutions are named GNN-based causal estimator (GNNCE), NetEst, and SPNet respectively. DDP stands for distribution discrepancy penalty.
We summarize the comparison of the past work on causal inference under networked interference using neural network and our work in Table 1. The most salient difference between our model and other methods (Ma and Tresp, 2021; Jiang and Sun, 2022; Huang et al., 2023) in this literature is that while the cited works deal with nonrandom treatment assignment, in our training dataset, students are randomly assigned to different classrooms conditional on observable school choice. Nevertheless, measuring the causal impact of peer effect is not straightforward as friendship network is not randomly assigned. Leveraging on this natural experiment setup of a conditional random classroom assignment, we construct valid instrumental variables to conduct causal inference analysis (Imbens and Angrist, 1995). See Appendix A.1 for a more detailed comparison between our work and this literature.
2.7 Genetic algorithm and algorithmic fairness
The question of optimal classroom assignment has been put forward but left unanswered in Berthelon et al. (2019). Bizzotto and Vigier (2022) and Sarkisian and Yamashita (2024) discuss this problem in a theoretical setup but do not bring it to empirical data. We adopt a genetic algorithm (GA) to solve this combinatorial optimization problem of class assignment with real data (Kumar et al., 2010; Alam et al., 2020). We further include a literature review on evolutionary algorithm in Appendix A.2. Despite GA’s potential benefits, there exists a risk of principals exploiting GA to prioritize maximizing the average peer effect at the expense of students in the lower quantiles. We contribute to the GA literature by considering algorithmic fairness, a concept extensively discussed by Kleinberg et al. (2018); Mitchell et al. (2021). When evaluating fitness in GA, we penalize the standard deviation of predicted peer effect both within and across classrooms. The modified GA termed algorithmically fair genetic algorithm (AFGA), outputs a more equitable class assignment policy. In addition, we demonstrate a trade-off between efficiency and equity when using GA versus AFGA.
2.8 Optimization methods for treatment/group assignment problems
Given the network formation prediction model (PeerNN) and peer effect estimated parameters, a key challenge in solving for the optimal class assignment policy is the dependency of the network structure on the policy itself. This sets our problem apart from the existing literature on treatment assignment under networked interference, which assumes a fixed network (See Appendix A.1). On top of the three papers cited in Section 2.6, there are numerous recent papers that look at the theoretical properties of methods that optimize treatment assignment given a fixed network (Ananth, 2020; Munro et al., 2021; Kitagawa and Wang, 2023a, b; Viviano, 2024). Another stream of literature that solves group assignment problems as either linear programming problems (LP) or mixed integer programming problems (MIP) assumes that link formation between any two individuals is independent of the presence of other individuals. This literature (Bhattacharya, 2009; Graham et al., 2014; Carrell et al., 2013) assumes a friendship formation process like the linear-in-means model, though Bhattacharya (2009); Graham et al. (2014) do allow heterogeneous peer effect parameters. The straightforward dependence between linkage formation and assignment policy allows the optimization problem to be recast as standard programming problems. On the other hand, our friendship formation model accommodates more sophisticated friendship formation patterns such as clustering or more general dependence between a pair of students’ friendship formation and other classmates’ characteristics. Such flexibility makes it impossible for us to recast our optimization problem as an LP or MIP. Due to the added complexity of our problem, we resort to a heuristic optimization method, namely, genetic algorithm (GA), instead of directly solving for an optimal policy, as in the cited works.
3 Data description and data notation

### v5

# Balancing Efficiency and Equity in Classroom Assignment under Endogenous Peer Effects

@@ Abstract
This paper presents a three-step empirical framework for optimizing classroom assignments under endogenous peer effects, using data from the China Education Panel Survey (CEPS).
Step 1: Modeling Friendship Networks. We design PeerNN, a neural network that mimics endogenous network formation as a discrete choice model, generating a friendship-intensity matrix ($\Omega$) that captures student popularity.
Step 2: Estimating Peer Effects. We measure the peer effect friends’ average 6th-grade class rank weighted by $\Omega$ on 8th-grade cognitive test score. Incorporating $\Omega$ into the linear-in-means model induces endogeneity. Using quasi-random classroom assignments, we instrument friends’ average 6th-grade class rank with the average classmates’ 6th-grade class rank (unweighted by $\Omega$). Our main regression result shows that a 10% improvement in friends’ 6th-grade class rank raises 8th-grade cognitive test scores by 0.13 SD. Positive $\beta$ implies maximizing (minimizing) the popularity of high (low) achievers optimizes outcomes.
Step 3: Simulating Policy Trade-offs. We use estimates from Step 1 and Step 2 to simulate optimal classroom assignments. We first implement a genetic algorithm (GA) to maximize average peer effect and observe a 1.9% improvement. However, serious inequity issues arise: low-achieving students are hurt the most in the pursuit of the higher average peer effect. We propose an Algorithmically Fair GA (AFGA), achieving a 1.2% gain while ensuring more equitable educational outcomes.
These results underscore that efficiency-focused classroom assignment policies can exacerbate inequality. We recommend incorporating fairness considerations when designing classroom assignment policies that account for endogenous spillovers.
@@ 1 Introduction
This paper proposes a three-step empirical framework for classroom assignment optimization in the presence of endogenous network formation and peer effect (Goulas and Monachou, 2024). Imagine a middle school has an incoming cohort of students who have just finished elementary school. The middle school principal needs to assign these students to classrooms. The principal has the following two beliefs and forms an assignment strategy based on these beliefs.
The first belief is endogenous network formation. Once the students are assigned to classrooms, they form friendship networks in their classrooms and receive more peer influence from their close friends. As a result, popular students exert more peer effects on the entire classroom. The second belief is heterogeneous peer quality, which leads to differential peer effects. The peer effect received by a student can be positive or negative depending on whether the friends of the said student are good students or disruptive peers. As a result of these two beliefs, the principal wants to strategically devise a classroom assignment policy that maximizes good students’ popularity and minimizes disruptive peers’ popularity, ultimately maximizing the average peer effect.
Correspondingly, this paper breaks the principal’s classroom assignment problem into three parts to output a classroom assignment policy that aligns with the beliefs and the principal’s strategy. Figure 1 summarizes the beliefs and strategy of the principal and their corresponding steps in our three-step empirical framework.
Figure 1: Beliefs and strategy of the principal (solid) and their corresponding steps (dashed).
To the best of our knowledge, our framework is the first empirical attempt for classroom assignment optimization under both endogenous network formation and peer effect. Given the novelty of the problem, we make some adaptations of current methods pulled in from different literature. Next, we briefly introduce each step’s general idea, related literature, and result summary.
Step 1: Modelling friendship formation
We predict middle school classroom friendship formation with students’ 6th-grade information as predictors. The goal of Step 1 is to figure out which students will be popular given a fixed set of classmates, as written in Figure 1.
@@
By combining two streams of network literature, we propose a new estimation method, named PeerNN, for linkage intensity. The first stream of literature predicts linkage using neural network. Mainstream link prediction studies use neural networks (Ahmed and Chen, 2016; Wu et al., 2016; Wang et al., 2014; Dhelim et al., 2022; Chunaev, 2020; Su et al., 2022; Samanta et al., 2021; Ling et al., 2023). Though both our work and these cited papers use neural networks to approximate complicated social network formation processes, we intentionally avoid techniques such as aggregating neighbors’ information used in the cited papers. We exclusively use students’ individual information as predictors. This approach is necessitated by the absence of linkage information under counterfactual classroom assignments. The second literature models friendship preferences using a micro-founded discrete choice (Graham, 2017; Mele, 2022). We incorporate the microfoundation concepts, such as discrete choice, homophily, and transitivity, into PeerNN’s architecture.
An advantage of PeerNN over many other network models is that it does not need expensive linkage data as a response variable. Instead, we can use easier-to-collect aggregate relational data(McCormick and Zheng, 2015; Breza et al., 2020; McCormick, 2020). ARD also imposes less of a threat to individual identification and other privacy-related problems.
Preview of Step 1 results: PeerNN captures social network characteristics such as gender homophily, outperforming the uniform friendship formation process implicitly assumed in traditional linear-in-means models widely used in peer effect literature (Guryan et al., 2009; Carrell and Hoekstra, 2010; Carrell et al., 2013; Burke and Sass, 2013).
Step 2: Measuring peer effect
We measure how friends’ average 6th-grade class rank affects students’ 8th-grade cognitive test scores. If the peer effect is positive, then students who did well in 6th-grade will be considered good students, and those who did poorly in 6th-grade will be considered disruptive peers.
@@
We replace the restrictive uniform spillover assumption of the linear-in-means model with a more reasonable friendship-weighted specification. This specification leverages the nonuniform friendship formation predicted by PeerNN, addressing the limitations of the linear-in-means model, which assumes uniform spillover effects and ignores friendship dynamics (Bramoullé et al., 2009; Goldsmith-Pinkham and Imbens, 2013). The uniform spillover assumption is especially problematic for our datasets, in which the classroom size is relatively large (Hong, 2022). The average classroom size is more than 40 students.
Introducing a network into pure contextual peer effect measurement is challenging because network formation is endogenous. Unobserved student characteristics can simultaneously affect friend choice and educational outcome. We exploit the random classroom assignment conditional on school choice to construct an instrument to conduct causal inference. We instrument friends’ average 6th-grade class rank (weighted by $\Omega$) with the classmates’ average 6th-grade class rank (unweighted by $\Omega$). This instrument satisfies the relevance constraint because the assigned classmates are the pool of friends that a student can pick from, so friends’ average and classmates’ average 6th-grade class rank correlate. The exclusion constraint is satisfied in theory because the classmates’ assignments are random after conditioning on school-fixed effects, assuming the absence of correlated peer effect mechanisms that do not operate through friendship. Such an assumption is likely to be violated, e.g., competitiveness of the classroom environment, quality of classroom discussion, and teachers’ adjustments for the level of instruction can correlate with the classroom assignment. To address these correlated mechanisms, we replace the classroom-random effect with classroom-level covariates to close the backdoor path. Furthermore, we test out the heterogeneous peer effect design with our dataset.
Preview of Step 2 results: All specifications’ conclusions are consistent: having friends who were high-achieving students in elementary school causally increases students’ 8th-grade cognitive test scores. In particular, the our baseline specification shows that a 10% improvement in friends’ 6th-grade class rank raises 8th-grade cognitive test scores by 0.13 SD.
Step 2 result means that students who did well in elementary schools are the good students whereas students who did poorly in elementary schools are the disruptive peers. This slightly differs from Figure 1 in the sense that instead of a binary classification, we have a spectrum of good students and disruptive peers since class ranks are converted into a continuous measure of class quantile.
Step 3: Optimizing classroom assignment
In the third step, we optimize classroom assignments in terms of maximizing average peer effects by maximizing/minimizing the popularity of the good students/disruptive peers, respectively.
Unlike existing literature, which implicitly assumes straightforward network formation that depends on only policy but not individuals’ characteristics (Bhattacharya, 2009; Carrell et al., 2013), our approach allows dependence between friendship intensity and characteristics of all classmates in one network, making the network dependent on both the assignment policy and the predetermined student characteristics (i.e., endogenous). Consequently, we cannot recast the optimization problem as a linear programming problem as done by the cited works. We employ a heuristic optimization method, genetic algorithm (GA), to navigate this complex problem.
@@
Preview of Step 3 results: GA outputs a classroom assignment policy that, at median, improves peer effect by 1.9% in our counterfactual simulation. However, the output policy is extremely inequitable. In pursuing the highest average peer effect for the whole classroom, GA gives up on a few disruptive peers by predicting that they form an exclusive clique. Consequently, students in that clique experience severely negative peer effects, but overall, the entire classroom has a higher average peer effect. A naive optimization of classroom assignment can lead to extremely inequitable educational outcomes. We propose a modified fairness-aware fitness function for GA and name it algorithmically fair genetic algorithm (AFGA) (the naming follows Kleinberg et al. (2018); Mitchell et al. (2021)). The new fitness function penalizes the variance of predicted peer effects within and across classrooms. AFGA outputs slightly less efficient (the improvement is 1.2%) but much more equitable policies.
Contribution of the paper
Our contributions are two-fold. First, we propose a novel way of estimating friendship formation parameters as a discrete choice model. The estimation method, named PeerNN, combines the flexibility of modern machine learning algorithms and the microfoundation of economic modeling. We apply PeerNN to CEPS data and show that it can capture prominent features in the network. Readers can use PeerNN as an exploratory tool to look for hidden network structures when aggregated relational data is available, or it can be used as a preliminary step for peer effect estimation, as in this paper. Second, though executing our framework in real life is challenging, our attempt to optimize classroom assignment highlights the critical need to incorporate fairness considerations when optimizing the efficiency of classroom assignment policies. In pursuit of optimal average outcome, students with weaker initial academic standing are hurt the most, causing inequity concerns. We recommend explicitly placing equity safeguards when designing optimal policy under an endogenous spillover setup, and propose an intuitive way to do so.
Paper organization
@@
The rest of the paper is organized as follows: Section 2 describes the CEPS data and introduces the notations that we use in this paper, Section 3 sets up the models and objective functions in our three-step framework, Section 4 demonstrates and explains the results that we obtain from the empirical study and Section 5 concludes the paper.
2 Notation and data description